OpenStudy (anonymous):

Suppose that a researcher approximates the density of the CAD/USD exchange rate by x~ N(1,0.0004) a) Find the probability that the exchange rate is less than 0.98

OpenStudy (slaaibak):

\[P(X< 0.98) = P({X - 1 \over \sqrt{0.0004}} <{ .98 - 1 \over \sqrt{0.0004}}) = P(Z < -1) = P(Z > 1)\] Now just use your normal tables.

OpenStudy (anonymous):

why do you have to divide by the square root of .0004? I thought that it was the standard deviation, not the variance?

OpenStudy (slaaibak):

Nope. The notation means X~N(mean, variance)

OpenStudy (slaaibak):

so taking the square root of .0004 gives us the standard deviation